OpenStudy (anonymous):

is it possible for a function to be both odd and even? all I can think of is the function { (0,0) }

OpenStudy (zarkon):

if f(-x)=f(x) and f(-x)=-f(x) then f(x)=-f(x) so 2f(x)=0 thusf(x)=0

OpenStudy (anonymous):

:O f(x) = 0 both odd and even